OpenStudy (anonymous):

5x + y = 6 2y = x +1

OpenStudy (anonymous):

Solve by substitution?

OpenStudy (anonymous):

5x + y=6. Solve for y is an option. Y=6-5x. Plug that into the 2nd eq & solve for x

OpenStudy (anonymous):

2(6-5x)= x+1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Once u get x, you can find y by substituting x back into either eq. Does that help?

OpenStudy (anonymous):

You will put your answer in an ordered pair. ( x,y)